GRE exam for university admission Test center 1 (if number of students is less than or equal to 100) (KD 250 Test center- extra room (if more than 100-in addition to the test center 1) Registration docs (KD 25 per student) Test center 2 fees (KD 200) test cost per student (KD 50)-AUK will pay the rest a. What is the total cost if 20 students registered for the test?
c. How many students do you need to breakeven?
Answer #1

a) If there are only 20 students, then we will need only test centre 1, which will cost KD 250

Cost of registration docs of 20 students = 25*20 = KD 500

Total cost if 20 students registered will be 250+500 = KD 750

c) At breakeven point , Total revenue = Total cost

If there are 100 students total revenue will be 50*100 = KD 5000 and total cost will be

= 250 + 25*100 = KD 2750

We see that total revenue exceed total cost , so total students will be less than 100. From this we know that we will not need test centre 2.

Let the number of students be x.

We will get the equation,

50 * x = 250 + 25 * x

=> 25x = 250

=> x = 10

So the breakeven point will be at 10 students. We can see that as total cost will be 250+ 25*10 = 500

and total revenue will be 50*10 = 500. This shows that total revenue is equal to total cost.